# Computational considerations for the prediction of airfoil Stall Flutter

### v1

@@ Abstract
This paper presents a comprehensive numerical investigation of a NACA0012 undergoing Stall Flutter Limit Cycle Oscillations (LCO) across distinct fluid dynamics regimes. It accurately models Small Amplitude Oscillations (SAO) in the transitional Reynolds regime and Large Amplitude Oscillations (LAO) in the moderate regime, observed in different experimental campaigns. The SAO analysis servs as a verification of the computational framework against established numerical benchmarks. Crucially, the LAO simulations represent the first documented prediction across the full experimental velocity range correlated against available measured data, addressing a significant literature gap. The predictions fidelity relies on rigorous computational criteria defined through a detailed sensitivity analysis. This demonstrated numerical requirements significantly more demanding than those typically employed for computing static polars or simulating dynamic pitching motion of rigid airfoils, underscoring the severity of the aeroelastic problem. Overall, the results show strong qualitative agreement with experimental observations, successfully reproducing key dynamic stall mechanics and bifurcation phenomena. Quantitatively, however, the simulation systematically over–predicts the critical onset velocity and under–predicts the LCO amplitudes, a discrepancy attributed to reduced aerodynamic excitation peaks.
@@ 1 Introduction
Stall flutter is a non–linear aeroelastic instability phenomenon resulting from the coupling between the flexibility of an aerodynamic structure (a wing) and the flow, in cases fundamentally driven by dynamic flow separation (dynamic stall). The resulting structural response ends up in a change in the pitch angle of the wing, which in turn may give rise to periodic complex flow separation and perhaps re–attachment on both sides of the lifting surface. This interaction between the fluid and the structure (Fluid Structure Interaction – FSI) creates self–excited oscillations and allows for the continued energy transfer from the free stream to the wing. Depending on the correlation between the aerodynamic energy input and the structural energy dissipation, the system’s dynamic behavior varies. Classical bending–torsion flutter involves the interaction between two (or more) structural modes, leading to self–excited vibrations of exponentially increasing amplitude that are mainly unrelated to flow separation (at least in the initial phase of the phenomenon) and usually occur in a linear flow regime (see Figure 1). On the other hand, stall flutter is strongly related to non–linear aerodynamic excitation phenomena arising in stall and post–stall angles of attack, leading in many cases to self–excited single degree of freedom (pitching) oscillations of limited amplitude that are usually referred to as Limit Cycle Oscillations (LCO). LCO represent a state of dynamic equilibrium where the time–averaged energy absorbed from the fluid over one cycle is balanced by the energy dissipated by structural damping. This instability is commonly encountered on similar aerodynamic shapes such as wings or rotating blades in many different applications, including helicopter rotor blades ham1966torsional, wind turbine blades ericsson2000role, turbomachinery blades mendelson1949aerodynamic, aircraft tails victory1943flutter and highly flexible wings noll2004investigation.
@@
Although initial experimental work on stall flutter dates back to the 1940s victory1943flutter, much of the related research has historically focused on the aerodynamic phenomenon of dynamic stall in the second half of the $20^{th}$ century halfman1951evaluation; ericsson1980dynamic; mccroskey1981phenomenon; rivera1992naca. Nevertheless, stall flutter is an aeroelastic phenomenon resulting in LCO introduced by complex and non–linear dynamic flow phenomena such as dynamic stall. Dynamic stall is a distinct unsteady aerodynamic phenomenon, in which an abrupt loss of aerodynamic loads occurs due to flow separation over a lifting surface undergoing unsteady motion that may not necessarily originate in aeroelastic responses. Yet, in stall flutter oscillations, dynamic stall serves as the main excitation factor that stabilizes the limited amplitude oscillations. Over recent years, various types of low–speed stall flutter oscillations have been observed experimentally, leading to several classifications, typically based on the dominant degree of freedom (e.g. pitch-dominated, plunge-dominated). Within the pitch-dominated classification, distinct behaviors are observed:
@@
  • Small Amplitude Oscillations (SAO) with symmetric pitch response in the range between $0^{\circ}$ and $10^{\circ}$ around $0^{\circ}$ mean value were first systematically observed and reported through wind tunnel experiments in the transitional Reynolds number regime $(4.5*10^{4}\leq Re_{c}\leq 1.3*10^{5})$ poirel2008self. This type of SAO arises because laminar boundary layer separation near the trailing edge region initiates the loss of dynamic stability at $0^{\circ}$ pitch angle, while the subsequent presence of a Laminar Separation Bubble (LSB) stabilizes the fluttering motion into LCO. This transitional instability regime has since been thoroughly investigated using various Computational Fluid Dynamics (CFD) tools. Early computational efforts to numerically investigate the flow patterns generated around the oscillating wing and the resulting aerodynamics loads utilised Large Eddy Simulations (LES) relying on imposed motion poirel2010aerodynamics, while later analyses successfully employed fully unsteady aeroelastic simulations based on a 2D Unsteady Reynolds Averaged Navier Stokes (URANS) approach, effectively reproducing the self–sustained pitch oscillations across the relevant Reynolds number range poirel2011computational; abbas2021aeroelastic. These numerical predictions are utilised in the present work for the verification of our comprehensive computational framework, allowing for crucial cross–reference against external data of comparable fidelity.
@@
  • Large Amplitude Oscillations (LAO) with symmetric pitch response in the range between $10^{\circ}$ and $40^{\circ}$ that are observed in a higher Reynolds number range $(2.5*10^{5}\leq Re_{c}\leq 5*10^{5})$ and are linked to the generation and shedding of a strong Leading Edge Vortex (LEV), similar to those observed during dynamic stall li2007experimental; dimitriadis2009bifurcation. In the same experimental campaign, asymmetric LCO of smaller amplitude (up to $20^{\circ}$) around the static equilibrium angle have been also observed at higher Reynolds numbers (up to $5*10^{5}\leq Re_{c}\leq 6*10^{5}$), resulting from the interaction between the stall flutter mechanism and the system’s static divergence. Contrary to SAO cases, accurate CFD predictions reproducing these large amplitude LCO experiments are relatively scarce in the literature yabili2012unsteady. At least to the authors’ knowledge, the successful computational replication of these large amplitude stall flutter LCO across the entire experimental Reynolds number range has not yet been thoroughly reported, highlighting a significant literature gap that this paper aims to address.
@@ 2 Experimental set–up
The following sections detail the experimental configurations utilised for the validation the computational framework presented in this paper, focusing on the distinct stall flutter regimes: Small Amplitude Oscillations (SAO) and Large Amplitude Oscillations (LAO). Sections and 4.2 provide specific details on the physical set–ups and structural parameters of these experiments, while Section 2.3, offers a general overview of the numerical approach.
@@ 2.1 Small Amplitude Oscillations
The validation case for SAO is based on experiments performed by Poirel, Harris and Benaissa poirel2008self in the closed circuit low–speed wind tunnel of the Royal Military College of Canada (RMC). This study focused on aeroelastic LCO occurring in the narrow transitional Reynolds number regime, specifically $4.5*10^{4}\leq Re_{c}\leq 1.3*10^{5}$. The model utilised a NACA0012 airfoil section (chord c=0.156mc=0.156\;m, span b=0.61mb=0.61\;m) that was structurally constrained to rotate in pure pitch, with plunging motion held fixed due to infinite stiffness, and was mounted with end–plates in order to minimize $3D$ effects. The apparatus possessed a mass moment of inertia $I_{s}=0.00135\;kgm^{2}$ and a structural stiffness $K_{s}=0.30\;Nm/rad$ and in the baseline configuration, the elastic axis (pitching axis) was located at $X_{ea}=0.18c$ aft of the leading edge. The system exhibited a viscous structural damping coefficient $D_{s}=0.002\;Nms$, resulting in a structural damping ratio $\zeta=0.05$ during free decay tests in still air. The self–sustained oscillations were typically observed for airspeeds between $5$ and 12m/s12\;m/s, corresponding to the transitional Reynolds range. The baseline value for the free–stream turbulence intensity of the wind tunnel was $T_{u}\approx 0.2\%$. These small amplitude oscillations ($\theta_{max}\leq 5.5^{\circ}$) were the first LCO associated with low Reynolds number effects to be systematically reported. The underlying mechanism is fundamentally linked to laminar boundary layer separation near the trailing edge, initiating the instability and leading to the designation laminar separation flutter. The motion was generally characterized by a well–behaved harmonic response and was measured with rotary potentiometers, whereas the flow periodicity was monitored using hot–wire anemometry in the wake. The aerodynamic moment was calculated as the right hand side of the dynamic equation of motion of the wing, by assuming a pure sinusoidal variation of the pitch angle of the wing with a known frequency and amplitude. A schematic of the experimental set–up is shown in Figure 2.
Figure 2: Schematic of the experimental set–up used to study Small Amplitude Stall Flutter Oscillations. Image adapted from poirel2008self.
@@ 2.2 Large Amplitude Oscillations
The LAO experiments used for validation were conducted at the University of Manchester by Dimitriadis and Li li2007experimental; dimitriadis2009bifurcation. Static measurements were carried out in the Project tunnel, which is an open–return subsonic wind tunnel with a working section of 2.16m2.16\;m in length, 1.1m1.1\;m in width and 0.86m0.86\;m in height, whereas dynamic measurements were conducted in the Avro $9x7\;ft^{2}$ closed–return subsonic wind tunnel. The test model featured a rectangular wing with a NACA0012 airfoil section, defined by a b=900mmb=900\;$m$ span and a c=300mmc=300\;$m$ chord. The support mechanism allowed for two degrees of freedom (pitch and plunge). However, the configuration was engineered to exhibit a single–mode dynamic response dominated by pitch motion due to the intentionally high plunge stiffness ($K_{h}=30.5\;N/mm$) compared to the pitch stiffness ($K_{p}=13.1\;Nm/rad$). The pitch axis was set $X_{ea}\approx 0.383c$ aft of the leading edge. The wing had a total weight of m=15kgm=15\;kg and a moment of inertia of $I_{p}=0.31\;kgm^{2}$ about the pitch axis. The system did not apply explicit damping mechanisms. Instead, the observed damping was primarily due to friction in the pitch springs, which led to non–linear structural damping. Analysis of free decay tests at 0, $2.5$ and 3.1m/s3.1\;m/s airspeed showed that the damping of the pitching mechanism was essentially linear at high pitch angles but became significantly non–linear at low pitch angles. The damping in plunge can be assumed to be very large. This campaign captured symmetric LCO with amplitudes reaching up to $40^{\circ}$, typically observed for airspeeds between $12$ and 25m/s25\;m/s that correspond to a moderate Reynolds number range extending from $2.4*10^{5}$ up to $5*10^{5}$. Furthermore, asymmetric LCO were observed at higher airspeeds between $25$ and 30m/s30\;m/s ($5*10^{5}\leq Re_{c}\leq 6*10^{5}$) resulting from the interaction of the flutter mechanism with the static divergence airspeed of the system. No self–sustained oscillations were observed for airspeeds between $3$ and 12m/s12\;m/s ($6*10^{4}\leq Re_{c}\leq 2.4*10^{5}$). The free–stream turbulence intensity in this large wind tunnel was found to be low, typically estimated to be around $T_{u}=0.05\%$. Unsteady pressure data were measured around the mid–span section, whereas the motion of the wing was measured using laser displacement probes. A picture of the experimental set–up is shown in Figure 3.
Figure 3: Image of the experimental set–up used to study Large Amplitude Stall Flutter Oscillations. Image copied from li2007experimental.
@@ 2.3 Computational set–up
Figure 4: Schematic of the computational set–up used to study the Small Amplitude Stall Flutter Oscillations measured in poirel2008self and the Large Amplitude Stall Flutter Oscillations measured in dimitriadis2009bifurcation. The aerodynamic analysis is performed through a 2D URANS CFD framework, whereas for the prediction of the system dynamics response, only the pitch DOF is considered.
@@ 3.1 The aerodynamic model
The core aerodynamic calculations for the aeroelastic simulations are performed using the in–house CFD solver MaPFlow papadakis2014development, which utilises a cell–centered finite volume discretisation scheme and is capable of handling both structured and unstructured computational grids. MaPFlow is built upon the compressible URANS equations described by Equation (1).
@@
In Equation (1), $\mathbf{U}$ is the vector of the Conservative Flow Variables:
@@
The convective fluxes $\mathbf{F_{c}}$, described by Equation (2), are evaluated using the Roe’s approximate Riemann solver roe1986characteristic described in Equation (3) coupled with the Venkatakrishnan limiter venkatakrishnan1993accuracy.
@@
In case Low Mach Preconditioning is used, the eigenvalues and the eigenvectors of the preconditioned system must be used ($\boldsymbol{\Lambda}_{\Gamma},\mathbf{R}_{\Gamma}$). Hence, $|\mathbf{A_{Roe}}|$ is changed to (according to Asouti_phd):
@@
MaPFlow can handle both steady and unsteady flows. The unsteady nature of FSI problems necessitates time–true computations using the implicit $2nd$ order backwards differences scheme described in Equation (6), combined with the dual time–stepping technique to facilitate convergence vatsa2010re.
@@
To ensure numerical stability and robustness in the low Mach number flow regime, Eriksson’s Preconditioning Matrix Eriksson1996 is consistently applied to the governing equations, due to its successful use in Asouti2005.

### v2

# Computational study of airfoil stall flutter Limit Cycle Oscillations

@@ Abstract
This paper presents a comprehensive numerical investigation of a NACA0012 undergoing Stall Flutter Limit Cycle Oscillations (LCO) across distinct fluid dynamics regimes. It accurately models Small Amplitude Oscillations (SAO) in the transitional Reynolds regime and Large Amplitude Oscillations (LAO) in the moderate regime, observed in different experimental campaigns. The SAO analysis serves as a verification of the computational framework against established numerical benchmarks. Crucially, the LAO simulations represent the first documented prediction across the full experimental velocity range correlated against available measured data, addressing a significant literature gap. The predictions fidelity relies on rigorous computational criteria defined through a detailed sensitivity analysis. This demonstrated numerical requirements significantly more demanding than those typically employed for computing static polars or simulating dynamic pitching motion of rigid airfoils, underscoring the severity of the aeroelastic problem. Quantitatively the simulation systematically over–predicts the critical onset velocity and under–predicts the LCO amplitudes.However, the results show strong qualitative agreement with experimental observations, successfully reproducing key dynamic stall mechanics and bifurcation phenomena.
@@ 1 Introduction
Stall flutter is a non–linear aeroelastic instability phenomenon resulting from the coupling between the flexibility of an aerodynamic structure (a wing) and the flow, in cases fundamentally driven by dynamic flow separation (dynamic stall). The resulting structural response ends up in a change in the pitch angle of the wing, which in turn may give rise to periodic complex flow separation and perhaps re–attachment on both sides of the lifting surface. This interaction between the fluid and the structure (Fluid Structure Interaction – FSI) creates self–excited oscillations and allows for the continued energy transfer from the free stream to the wing and vice–versa. Depending on the correlation between the aerodynamic energy input and the structural energy dissipation, the system’s dynamic behavior varies. Classical bending–torsion flutter involves the interaction between two (or more) structural modes, leading to self–excited vibrations of exponentially increasing amplitude that are mainly unrelated to flow separation (at least in the initial phase of the phenomenon) and usually occur in a linear flow regime (see Figure 1). On the other hand, stall flutter is strongly related to non–linear aerodynamic excitation phenomena arising in stall and post–stall angles of attack, leading in many cases to self–excited single degree of freedom (pitching) oscillations of limited amplitude that are usually referred to as Limit Cycle Oscillations (LCO). LCO represent a state of dynamic equilibrium where the time–averaged energy absorbed from the fluid over one cycle is balanced by the energy dissipated by structural damping. This phenomenon is typically a single degree of freedom instability, that only occurs if the pitching eigen–frequency is lower than the plunging one, or if plunging is restricted. It is commonly encountered on similar aerodynamic shapes such as wings or rotating blades in many different applications, including helicopter rotor blades Ham and Young (1966), wind turbine blades Ericsson (2000), turbomachinery blades Mendelson (1949), aircraft tails Victory (1943) and highly flexible wings Noll et al. (2004).
@@
Although initial experimental work on stall flutter dates back to the 1940s Victory (1943), much of the related research has historically focused on the aerodynamic phenomenon of dynamic stall in the second half of the $20^{th}$ century Halfman et al. (1951); Ericsson and Reding (1980); McCroskey (1981); RIVERA et al. (1992). Nevertheless, stall flutter is an aeroelastic phenomenon resulting in LCO introduced by complex and non–linear dynamic flow phenomena such as dynamic stall. Dynamic stall is a distinct unsteady aerodynamic phenomenon, in which an abrupt loss of aerodynamic loads occurs due to flow separation over a lifting surface undergoing unsteady motion that may not necessarily originate in aeroelastic responses. Yet, in stall flutter oscillations, dynamic stall serves as the main excitation factor that stabilizes the limited amplitude oscillations. Over recent years, various types of low–speed stall flutter oscillations have been observed experimentally, leading to several classifications, typically based on the dominant degree of freedom (e.g. pitch-dominated, plunge-dominated). Within the pitch-dominated classification, distinct behaviors are observed:
  • Small Amplitude Oscillations (SAO)
  • Large Amplitude Oscillations (LAO)
@@
SAO with symmetric pitch response in the range between $0^{\circ}$ and $10^{\circ}$ around $0^{\circ}$ mean value were first systematically observed and reported through wind tunnel experiments by Poirel et al. (2008) in the transitional Reynolds number regime $(4.5*10^{4}\leq Re_{c}\leq 1.3*10^{5})$. The loss of the dynamic stability at $0^{\circ}$ pitch angle was initially attributed to the laminar boundary separation near the trailing edge region, while the subsequent presence of a Laminar Separation Bubble (LSB) was considered as a stabilisation factor of the fluttering motion into LCO. The presence of a LSB is computationally verified in Poirel and Yuan (2010) and in the present study (see Section 4.1). Nevertheless, simulations indicate a Kelvin–Helmholtz instability occurring in the shear layer close to the trailing edge as the onset mechanism to the fluttering motion and the resulting von Kármán vortex shedding as the LCO stabilisation factor. This transitional instability regime has since been thoroughly investigated using various Computational Fluid Dynamics (CFD) tools. Early computational efforts to numerically investigate the flow patterns generated around the oscillating wing and the resulting aerodynamics loads utilised Large Eddy Simulations (LES) relying on imposed motion Poirel and Yuan (2010), while later analyses successfully employed fully unsteady aeroelastic simulations based on a 2D Unsteady Reynolds Averaged Navier Stokes (URANS) approach, effectively reproducing the self–sustained pitch oscillations across the relevant Reynolds number range Poirel et al. (2011); Abbas et al. (2021). These numerical predictions are utilised in the present work for the verification of our comprehensive computational framework, allowing for comparison against external data of comparable fidelity.
@@
LAO with symmetric pitch response in the range between $10^{\circ}$ and $40^{\circ}$ were later observed by Li (2007); Dimitriadis and Li (2009) in a higher Reynolds number range $(2.5*10^{5}\leq Re_{c}\leq 5*10^{5})$ and were linked to the generation and shedding of a strong Leading Edge Vortex (LEV), similar to those observed during dynamic stall. In the same experimental campaign, asymmetric LCO of smaller amplitude (up to $20^{\circ}$) around the static equilibrium angle have been also observed at higher Reynolds numbers (up to $5*10^{5}\leq Re_{c}\leq 6*10^{5}$), resulting from the interaction between the stall flutter mechanism and the system’s static divergence. Contrary to SAO cases, accurate CFD predictions reproducing these large amplitude LCO experiments are scarce in the literature. At least to the authors’ knowledge, the successful computational replication of these large amplitude stall flutter LCO across the entire experimental Reynolds number range has not yet been thoroughly reported, highlighting a significant literature gap that this paper aims to address.
@@ 2 Experimental set–up
The following sections detail the experimental configurations utilised for the validation the computational framework presented in this paper, focusing on the distinct stall flutter regimes: Small Amplitude Oscillations (SAO) and Large Amplitude Oscillations (LAO). Sections and 2.2 provide specific details on the physical set–ups and structural parameters of these experiments, while Section 2.3, offers a general overview of the numerical approach.
@@ 2.1 Small Amplitude Oscillations
The validation case for SAO is based on experiments performed by Poirel, Harris and Benaissa Poirel et al. (2008). This study focused on aeroelastic LCO occurring in the narrow transitional Reynolds number regime, specifically $4.5*10^{4}\leq Re_{c}\leq 1.3*10^{5}$. The model utilised a NACA0012 airfoil section (chord c=0.156mc=0.156\;m, span b=0.61mb=0.61\;m) that was structurally constrained to rotate in pure pitch, with plunging motion held fixed due to infinite stiffness, and was mounted with end–plates in order to minimize $3D$ effects. The apparatus possessed a mass moment of inertia $I_{s}=0.00135\;kgm^{2}$ and a structural stiffness $K_{s}=0.30\;Nm/rad$ and in the baseline configuration, the elastic axis (pitching axis) was located at $X_{ea}=0.18c$ aft of the leading edge. The system exhibited a viscous structural damping coefficient $D_{s}=0.002\;Nms$, resulting in a structural damping ratio $\zeta=0.05$ during free decay tests in still air. The self–sustained oscillations were typically observed for airspeeds between $5$ and 12m/s12\;m/s, corresponding to the transitional Reynolds range. The baseline value for the free–stream turbulence intensity of the wind tunnel was $T_{u}\approx 0.2\%$. These small amplitude oscillations ($\theta_{max}\leq 5.5^{\circ}$) were the first LCO associated with low Reynolds number effects to be systematically reported. The underlying mechanism is fundamentally linked to laminar boundary layer separation near the trailing edge, initiating the instability and leading to the designation laminar separation flutter. The motion was generally characterized by a well–behaved harmonic response and was measured with rotary potentiometers, whereas the flow periodicity was monitored using hot–wire anemometry in the wake. The aerodynamic moment was calculated as the right hand side of the dynamic equation of motion of the wing, by assuming a pure sinusoidal variation of the pitch angle of the wing with a known frequency and amplitude. A schematic of the experimental set–up is shown in Figure 2.
Figure 2: Schematic of the experimental set–up used to study Small Amplitude Stall Flutter Oscillations. Image adapted from Poirel et al. (2008).
@@ 2.2 Large Amplitude Oscillations
The LAO experiments used for validation were conducted at the University of Manchester by Dimitriadis and Li Li (2007); Dimitriadis and Li (2009). Static measurements were carried out in the Project tunnel, which is an open–return subsonic wind tunnel with a working section of 2.16m2.16\;m in length, 1.1m1.1\;m in width and 0.86m0.86\;m in height, whereas dynamic measurements were conducted in the Avro $9x7\;ft^{2}$ closed–return subsonic wind tunnel. The test model featured a rectangular wing with a NACA0012 airfoil section, defined by a b=900mmb=900\;$m$ span and a c=300mmc=300\;$m$ chord. The support mechanism allowed for two degrees of freedom (pitch and plunge). However, the configuration was engineered to exhibit a single–mode dynamic response dominated by pitch motion due to the intentionally high plunge stiffness ($K_{h}=30.5\;N/mm$) compared to the pitch stiffness ($K_{p}=13.1\;Nm/rad$). The pitch axis was set $X_{ea}\approx 0.383c$ aft of the leading edge. The wing had a total weight of m=15kgm=15\;kg and a moment of inertia of $I_{p}=0.31\;kgm^{2}$ about the pitch axis. The system did not apply explicit damping mechanisms. Instead, the observed damping was primarily due to friction in the pitch springs, which led to non–linear structural damping. Analysis of free decay tests at 0, $2.5$ and 3.1m/s3.1\;m/s airspeed showed that the damping of the pitching mechanism was essentially linear at high pitch angles but became significantly non–linear at low pitch angles. The damping in plunge can be assumed to be very large. This campaign captured symmetric LCO with amplitudes reaching up to $40^{\circ}$, typically observed for airspeeds between $12$ and 25m/s25\;m/s that correspond to a moderate Reynolds number range extending from $2.4*10^{5}$ up to $5*10^{5}$. Furthermore, asymmetric LCO were observed at higher airspeeds between $25$ and 30m/s30\;m/s ($5*10^{5}\leq Re_{c}\leq 6*10^{5}$) resulting from the interaction of the flutter mechanism with the static divergence airspeed of the system. No self–sustained oscillations were observed for airspeeds between $3$ and 12m/s12\;m/s ($6*10^{4}\leq Re_{c}\leq 2.4*10^{5}$). The free–stream turbulence intensity in this large wind tunnel was found to be low, typically estimated to be around $T_{u}=0.05\%$. Unsteady pressure data were measured around the mid–span section, whereas the motion of the wing was measured using laser displacement probes. A picture of the experimental set–up is shown in Figure 3.
Figure 3: Image of the experimental set–up used to study Large Amplitude Stall Flutter Oscillations. Image copied from Li (2007).
@@ 2.3 Computational set–up
Figure 4: Schematic of the computational set–up used to study the Small Amplitude Stall Flutter Oscillations measured in Poirel et al. (2008) and the Large Amplitude Stall Flutter Oscillations measured in Dimitriadis and Li (2009). The aerodynamic analysis is performed through a 2D URANS CFD framework, whereas for the prediction of the system dynamics response, only the pitch DOF is considered.
@@ 3.1 The aerodynamic model
The core aerodynamic calculations for the aeroelastic simulations are performed using the in–house CFD solver MaPFlow Papadakis (2014b), which utilises a cell–centered finite volume discretisation scheme and is capable of handling both structured and unstructured computational grids. MaPFlow is built upon the compressible URANS equations described by Equation (1).
@@
In Equation (1), $\vec{U}$ is the vector of the Conservative Flow Variables:
@@
The convective fluxes $\mathbf{F_{c}}$, described by Equation (2), are evaluated using the Roe’s approximate Riemann solver Roe (1986) described in Equation (3) coupled with the Venkatakrishnan limiter Venkatakrishnan (1993).
@@
In case Low Mach Preconditioning is used, the eigenvalues and the eigenvectors of the preconditioned system must be used ($\boldsymbol{\Lambda}_{\Gamma},\mathbf{R}_{\Gamma}$). Hence, $|\mathbf{A_{Roe}}|$ is changed to (according to Asouti (2009)):
@@
MaPFlow can handle both steady and unsteady flows. The unsteady nature of FSI problems necessitates time–true computations using the implicit $2nd$ order backwards differences scheme described in Equation (6), combined with the dual time–stepping technique to facilitate convergence Vatsa et al. (2010).
@@
To ensure numerical stability and robustness in the low Mach number flow regime, Eriksson’s Preconditioning Matrix Eriksson (1996) is consistently applied to the governing equations, due to its successful use in Asouti et al. (2005).